# Task and Motion Planning in Hierarchical 3D Scene Graphs

### v1

@@ II-C Inferring the Planning Domain from Scene Graphs
As a running example for the remainder of this paper, we define an example problem using these predicates, motivated by CBRNE scenarios. In this “Inspection Domain”, an agent can be commanded to visit or avoid certain places, and inspect and neutralise objects that have been marked as suspicious. These suspicious objects cannot be passed until they have been inspected and neutralised. We therefore define problem specific predicates: (VisitedPlace ?p) which indicate the current and past places the robot has been to, and (Safe ?o) or (Suspicious ?o) which describe an object. We define sub-problems for planning motion between two poses, sampling poses for inspecting and neutralizing objects, and sampling poses in a specific place. Goal specifications in this domain can include positive or negated facts based on these predicates. The agent’s available actions are to move between poses in connected places, and to inspect objects from appropriate poses (for simplicity we do not separate the inspect and neutralise actions). Any valid plan is composed of these actions, which at execution time are converted into a motion sequence of FollowPath($t_{i}$) and InspectObject($o_{i}$) primitives for paths $t_{i}$ and objects $o_{i}$. Places and objects are the most numerous nodes in the scene graph, so we need to ensure solving tasks in our domain remains tractable as the number of places and objects grows. The complexity of dealing with places and objects will have the largest implication for the move action, which we present here:
  (:action move
    :parameters (?p1 ?p2 ?c1 ?c2 ?t)
    :precondition (and
              (Trajectory ?c1 ?t ?c2)
              (PoseInPlace ?c1 ?p1)
              (PoseInPlace ?c2 ?p2)
              (Connected ?p1 ?p2)
              (AtPose ?c1)
    :effect (and (AtPose ?c2)
                 (not (AtPose ?c1))
                 (VisitedPose ?c2))))
  (:derived (AtPlace ?p)
    (exists (?pose)
            (and (AtPose ?pose)
                 (PoseInPlace ?pose ?p))))
This action allows the agent to move between its current place $p_{1}$ and any place $p_{2}$ that is connected to $p_{1}$ in the scene graph, and updates the state to note the agent’s current pose and that $p_{2}$ has been visited. We refer to this problem representation as the direct encoding of the Inspection domain.
The direct encoding of the problem has one major drawback: moving between places that are not physically near each other necessitates chaining together potentially many move actions, which requires motion planning in many short segments between neighboring places. For simple navigation, planning in short segments can actually aid in the efficiency of long-distance motion planning, as we will discuss in Sec.III-C. However, relying on a general task planner to construct long-horizon action sequences, any segment of which may be invalidated by the low-level geometry, leads to very inefficient planning in practice. In the Inspection domain for example, solving a motion-planning sub-problem for a trajectory requires accounting for Suspicious objects that may block the way. However, there is no a priori logical connection in the PDDL problem between a trajectory and the objects which might interfere with it. Therefore, choosing the right object to inspect in order to unblock a path is difficult for the task planner. The longer the sequence of move actions needed to reach a goal state, the more potential plans the solver must consider. This scales poorly both in the size of the environment, and as we consider increasingly complex goal specifications (Sec. IV).
III Scalable Scene Graph Planning
Our objective is to both enable, and reduce the computational burden of solving TAMP problems in large environments. We emphasize three main criteria in our development of a DSG planning framework:
  1. 1.
    Generating plans that we know can be executed.
  2. 2.
    Preserving the set of goals that can be specified and solved.
  3. 3.
    Planning at large scale.
Solving problem instances with the direct encoding described above is computationally intractable for large scenes. To address this potential intractability, we first define the characteristics of a planning domain that, if met, would enable sparsification while still meeting the criteria listed above. Next, we propose a method for constructing a simplified planning domain that has these characteristics and enables planning in large scenes. Critically, our method depends on factoring the task and motion planning problem into separate task, navigation, and motion planning problems, aligned with different layers of the scene graph hierarchy. Finally, we describe our approach for planning, using feedback from failed attempts to solve sub-problems to add facts to our domain.
III-A Simplifying Problem Instances
One method for reducing the complexity of a planning problem is to shorten the depth and branching factor of the search that produces a plan. In the case of the direct encoding of our Inspect Domain, one way to do this would be to prune Place symbols from the planning domain that are not relevant to the problem instance, and redefine the move action to allow for travel over longer distances. This pruning would remove possible regions the robot might travel and potentially reduce the depth of search required to find a plan, thus simplifying the search problem. However, identifying symbols that do not impact the solution is in general as hard as solving the planning problem itself, and simply removing places from a problem instance might render the problem as encoded in Section II-C infeasible if it is done naively. In this section, we will show that slight modifications to the domain and motion planning sub-problem lead to straightforward identification of symbols that can safely be removed from the planning problem. We begin by defining a set of symbols that are redundant for a particular goal specification.
@@ III-A Simplifying Problem Instances
The brief intuition behind this notion of redundancy is that 1) if any plan involving the symbol yields a motion sequence that can be rewritten without the symbol, the symbol is redundant, and 2) if we solve a planning instance where a redundant symbol has been removed, we would like to know that the plan is still valid in the original problem. The requirements for a symbol to be redundant are quite strong (every plan that uses a symbol must have an alternate plan that does not use the symbol and still results in the same motion sequence), but we will later show that a slight modification to the movement action lets many places have this property. We now prove that we do not lose anything by removing a redundant symbol.
@@
Consider a feasible planning instance $(\mathcal{P},\mathcal{A},\mathcal{S},\mathcal{O},\mathcal{I}_{0},\mathcal{G})$. For a redundant symbol $x\in\mathcal{O}$, we define a related instance $R^{\prime}=(\mathcal{P},\mathcal{A},\mathcal{S},\mathcal{O}^{\prime},\mathcal{I}_{0}^{\prime},\mathcal{G}^{\prime})$ where $x$ has been removed, i.e., $\mathcal{O}^{\prime}=\mathcal{O}\setminus x$ and $\mathcal{I}_{0}^{\prime}$ contains all facts in $\mathcal{I}_{0}$ except those parameterized by $x$, and similarly for $\mathcal{G}^{\prime}$. Let $\Pi_{R}$ denote the set of valid plans for $R$. Then, $\Pi_{R^{\prime}}\subseteq\Pi_{R}$ and $\Pi_{R^{\prime}}\neq\emptyset$.
Proof.
Consider $\pi\in\Pi_{R}$. If $\pi$ does not contain any actions parameterized by $x$, then the same plan $\pi$ is also a valid solution for $R^{\prime}$. Consider the alternative case, where $\pi$ does contain an action parameterized by $x$. By Definition 1, there is another plan $\pi^{\prime}$ with equivalent motion sequence not parameterized by $x$, which is a valid solution for $R^{\prime}$.
Now that we have shown that $\Pi_{R^{\prime}}$ is not empty, we need to show that any valid plan for $R^{\prime}$ is valid for $R$. Consider plan $\pi=[a_{1},...,a_{N}]\in\Pi_{R^{\prime}}$ with corresponding state plan $\mathcal{I}_{\pi}=[\mathcal{I}_{0},...,\mathcal{I}_{N}]$. If the addition of facts $\mathcal{F}$ parameterized by $x$ make $\pi$ invalid, then there must exist a state $\mathcal{I}_{k}$ such that $\mathcal{I}_{k}\cup\mathcal{F}\notin\text{Pre}(a_{k+1})$, which means that $a_{k+1}$ is parameterized by a symbol that did not exist in $R^{\prime}$. Only a universal or existential quantifier in $\text{Pre}(a_{k+1})$ can cause $a_{k+1}$ to be parameterized by an additional symbol. Adding additional facts cannot turn an existentially quantified formula from true to false. By Definition 1, $a_{k+1}$ does not have any universal quantifiers that can be parameterized by $x$ in its precondition. Thus $\pi$ must be valid for $R$.
In the original direct domain encoding presented in Sec. II-C, there are no redundant places, since removing any place implies that no other plans could result in a trajectory going through that place, which may be necessary depending on the geometry of the scene. However, let us consider a domain with a more general move action: moveRelaxed. This action takes $N$ places as parameters and has a precondition that a sub-problem is solved for a trajectory that goes through each of these $N$ places (and possibly others not included in the set of $N$). Trajectory represents this trajectory.
  (:action moveRelaxed
    :parameters (?p1 ... ?pN ?c1 ?t ?c2)
    :precondition (and
        (Trajectory ?p1 ... ?pN ?c1 ?c2 ?t)
        (AtPose ?c1)
    :effect (and (AtPose ?c2)
                 (not (AtPose ?c1))
                 (VisitedPlace ?p1)
                 (VisitedPlace ?pN)
                 (VisitedPose ?c2))))
With these modifications, potentially many places become redundant. For any place that appears in a plan incidentally (i.e., not part of the goal), there now exists an equivalent plan where these intermediate places do not appear.
@@
For example, imagine our robot begins in Place A, and is tasked with (among other things) inspecting an object in place C. A motion sequence corresponding to a plan to move from Place A to Place B, then from Place B to Place C can be equivalent to a sequence generated from a plan to move from Place A to C directly, so it seems that we can remove Place B from the problem. Extrapolating this pruning approach to larger scenes and more complex goals has the potential to vastly reduce planning horizons. However, now consider that the robot was also instructed to avoid Place B. The ability to include a constraint on the goal states of the form (not (VisitedPlace B)) complicates the pruning process. Executing moveRelaxed from Place A to Place C may involve following a trajectory that takes the robot through Place B, even if the goal specifies that Place B should not be visited, whether or not Place B is in problem domain. Technically this is still a valid solution to the planning problem since place B never appears as a parameter to the moveRelaxed action (and therefore (VisitedPlace B) is not an effect), but clearly the domain with a relaxed movement action does not fully capture how the user wants the robot to interact with the environment.
III-B Execution-Consistency
@@
To formalize the discrepancy between what happens when the robot executes a motion sequence and the constraints that we expect a planning problem to impose, we introduce the concept of a verifier function. A verifier function maps motion sub-sequences to sets of PDDL domain facts, and “verifies” which additional domain facts must be implicitly true as a result of the agent executing a motion sequence, even if actually adding these facts to the problem instance during the solving process is undesirable computationally. For example, if we want to consider if an agent’s trajectory avoids certain places, we can imagine a verifier that would return (VisitedPlace $p_{i}$) for each place $p_{i}$ that the continuous trajectory overlaps with. Given a verifier $V$, the facts that hold at each step when executing a motion sequence may be different than expected in the original plan. We denote the facts that would be added by such a verifier applied to the motion sequence associated with $a_{i}$ as $V(a_{i})$, and term this sequence of expanded states the $V$-extended state plan.
@@
For an action plan $\pi=[a_{1},...,a_{n}]$, its corresponding state plan $\mathcal{I}_{\pi}=[\mathcal{I}_{0},...,\mathcal{I}_{n}]$, and hypothetical verifier function $V$, the V-extended state plan for state $\mathcal{I}_{k}$ is:
$\mathcal{I}^{\prime}_{1}=\mathcal{I}_{1}\cup V(a_{1})$
@@
An extended state $\mathcal{I}^{\prime}_{k}$ is composed of the facts $\mathcal{I}_{k}$ in the initial plan, plus any extra facts that were present in the previous extended state $\mathcal{I}^{\prime}_{k-1}$ other than those removed by action $a_{k}$, plus any facts that would be returned by a verifier applied to action $a_{k}$. Informally, $\mathcal{I}^{\prime}_{k}$ is the state at step $k$ as experienced by the verifier. As discussed in Sec. II-A, any state plan found by a search algorithm is valid by construction. However, a state plan that is augmented with the extra facts that would be produced by a verifier might not be, for example, if the goal specification has a condition that the robot should avoid certain places in the domain.
To address this problem, let us imagine a verifier $V_{place}$ that takes a motion sub-sequence $\mu$, and returns a VisitedPlace fact for each place that intersects with the agent’s position while executing $\mu$. For a place $p$ and a trajectory $t$ to be followed by the motion primitive $\texttt{FollowPath}(t)$, we denote $p\cap t$ the section of $t$ that intersects with $p$. We can then define the verifier as
$\begin{split}V_{place}(\mu)=\{&\texttt{(VisitedPlace p)}~{}|~{}\\ &p\cap t_{i}\neq\emptyset\text{ for }\texttt{FollowPath}(t_{i})\in\mu\}.\end{split}$ (1)
@@
If the motion sequence associated with the action plan would result in the agent visiting a place that we do not expect, then the $V_{place}$-extended state plan would include a VisitedPlace fact that may conflict with the goal. From this idea, we define the concept of execution-consistency, which requires that solutions to the planning problem are still valid after considering the facts from a verifier.
Definition 3 (Execution-consistent).
@@
A domain is execution-consistent with respect to verifier $V$ if, for every valid plan $\pi$, the V-extended state plan is valid.
@@
A domain is trivially execution-consistent for the empty verifier $V(\cdot)=\emptyset$, as the extended state plan is equal to the original plan. A domain is also execution-consistent if the range of $V$ applied to each motion subsequence $\mu_{i}$ corresponding to action $a_{i}$ is limited to facts in $\text{Eff}(a_{i})$. In other cases, a domain can still be execution-consistent for a verifier that would introduce new facts if the domain is carefully crafted. In defining a planning domain for any task, we seek to have it execution-consistent with respect to any defined verifiers.

### v2

@@ 2.3 Inferring the Planning Domain from Scene Graphs
As a running example, we define an example problem using these predicates, motivated by CBRNE scenarios. In our “Inspection Domain”, an agent can be commanded to visit or avoid certain places, and inspect and neutralize objects that have been marked as suspicious. The robot cannot move past a suspicious object until it has been inspected and neutralized. We therefore define problem specific predicates: (VisitedPlace ?p) which indicates the current and past places the robot has been to, and (Safe ?o) or (Suspicious ?o) which describe an object. We define streams for sampling poses for inspecting and neutralizing objects, sampling poses in a specific place, and planning motion between two poses in order to find feasible continuous parameters for a given abstract plan. Goal specifications in this domain can include positive or negated facts based on these predicates. The agent’s available actions are to move between poses in connected places, and to inspect objects from appropriate poses (for simplicity we do not separate the inspect and neutralize actions). Note that only move needs to be parameterized by a place symbol. Any valid plan is composed of these actions, which at execution time are converted into a motion sequence of FollowPath($t_{i}$) and InspectObject($o_{i}$) primitives for paths $t_{i}$ and objects $o_{i}$.
3 Scalable Scene Graph Planning
Our objective is to both enable and accelerate solving TAMP problems in large environments. One associated challenge is that the number of symbols created by a scene graph can quickly overwhelm the ability of the planner to reason efficiently due to increasing branching factor, and the depth of search needed to find plans. However, we notice that the vast majority of planning domains, and the world in general, tend to factor into sequence of navigation actions punctuated by periodic object-centric actions. This factorization allows us to identify a subset of symbols relevant for object interaction, and a subset needed to move from place to place, potentially simplifying search. We therefore propose a planning formulation that aligns with the scene graph hierarchy and naturally divides the planning problem into a high-level, task-relevant planning problem such as finding a sequence of manipulation actions, mid-level coarse navigation planning between locations, and low-level continuous trajectory planning between points.
Specifically, instead of requiring a PDDL planner to find paths through the places in the scene graph at the discrete symbolic level, we reduce the depth of the planning horizon at the highest level by reasoning only over places that are directly relevant for achieving the goal. Then, a coarse navigation planner plans through the 2D places layer to create an abstract motion plan composed of a sequence of subgoals. Finally, a fine-grained motion planner is guided by the navigation plan subgoals through the places layer, quickly finding motion plans over large distances. We discuss this in Sec. 3.1.
By focusing each layer of our tri-level planner hierarchy on specific types of actions, we can prune irrelevant symbols within each layer and simplify the corresponding problems by reducing the branching factor of search (Sec. 3.2). Critical to our approach is that the proposed factorization must not limit the types of problems that can be solved, nor produce plans which violate intended constraints. To that end, we show a sufficient condition for symbols to be removed from each planning problem while maintaining feasibility, then show how to reason about whether the resulting motion plans adhere to the original specification (Sec. 3.3). Finally, in Sec 3.4, we consider an additional heuristic that enables optimistically ignoring objects that are irrelevant due to scene geometry.
3.1 Hierarchical Planning
Figure 2: Our tri-level planner in a real world environment. A scene graph of the fifth floor of building 45 at MIT was built using the Spot robot, from which we extract our planning abstraction for the goal: (and (ObjectAtPlace O27 P909) (VisitedPlace P2700) (Safe O35) (not (VisitedPlace P1153))), which instructs the agent to move O27 to a P909, inspect O35, visit P2700, all while avoiding P1153. At the highest level, the task planner is given a very sparsified version of the scene, as highlighted above. The mid-level planner plans a path through the places guided by the abstract plan found at the highest level, avoiding place P1153. Feedback from this level leads to the addition of O28 to the high-level domain, as O27 would be otherwise unreachable. The low-level planner computes full trajectories, guided by the path found at the mid-level. The plan produced (and executed by the robot) is shown on the right.
Here we describe our tri-level planning approach. The highest-level of the planner must reason about which objects to interact with, which regions to avoid, and which destinations the robot should move to. The most straightforward encoding of planning motion through a scene graph would closely mirror the connectivity of the graph, modeling the move action as a transition between two places that share an edge in the graph. We refer to this problem encoding as the direct encoding of the domain. However, as the scale of the environment increases, this direct encoding results in very long horizon plans that are expensive for the general-purpose PDDL planner to find.
Instead, we propose a more general move action: moveRelaxed. This action takes place $p_{1}$ and $p_{2}$ as parameters, in addition to initial and final poses $c_{1}$ and $c_{2}$ and a trajectory symbol $t$. The action’s effect moves the robot’s pose from $c_{1}$ to $c_{2}$, and marks $p_{1}$ and $p_{2}$ as visited. For example, consider the task in Fig. 1B where the robot begins in Place $P_{0}$, and is tasked with visiting place $P_{5}$. A motion sequence corresponding to a plan to move from place $P_{0}$ to place $P_{1}$, then from $P_{1}$ to $P_{4}$, and finally from $P_{4}$ to $P_{5}$ can be equivalent to a sequence generated from a plan to move from $P_{0}$ to $P_{5}$ directly, meaning the high-level planner need not explicitly plan to move through $P_{1}$ and $P_{4}$. Extrapolating this pruning approach to larger scenes and more complex goals has the potential to vastly reduce planning horizons, although it imposes certain constraints on the lower-level planners that will be addressed at length in Section 3.3.
Abstract plans produced by the high-level planner do not initially contain information about how the robot moves from the start to end poses. Instead, they optimistically contain trajectory symbols with continuous parameters that must be filled in by the lower-level planners. In order to do this efficiently, we rely on the 3D scene graph to accelerate motion planning. To find a motion plan between two configurations $c_{1}$ and $c_{2}$, we plan through the places layer of the scene graph, finding a sequence of places that leads from $c_{1}$ to $c_{2}$ and respects the connectivity of the scene graph. This level of abstraction can take advantage of Euclidean distance heuristics to accelerate planning, while allowing for the constraints of the task (e.g., avoiding a particular place) to be encoded simply.
At the lowest level of abstraction, the planner generates a kinematically-feasible path for the robot to follow based on the reference path from the mid-level planner. This path can be generated efficiently by first considering an optimistic path that connects the waypoints on the reference path and ignores obstacles. Any segments of this path that are rendered infeasible by obstacles or violations of kinematic constraints can be re-solved by a planner that considers obstacles, such as RRT [19]. Better alignment between which edges are present in the scene graph and kinematic feasibility for the robot leads to better performance of this heuristic. Fig. 2 highlights our tri-level planner in the real world.
3.2 Removing Redundant Symbols
Our relaxed encoding reduces the planning depth significantly for the high-level planner, but it introduces a different problem of branching factor. Now the high-level planner can consider moving between any two places as a valid action, which can make search difficult in certain problem instances. To address this concern, we can reduce the size of the planning instance by pruning Place symbols that are not relevant to a given planning problem. However, identifying symbols that do not impact the solution is in general as hard as solving the problem itself, and naively removing places from a problem instance might render the problem infeasible. In this section, we characterize a set of places that we know can safely be removed from the problem before planning given the semantics of moveRelaxed. We begin by defining a set of symbols that are redundant for a particular goal specification.
@@ 3.2 Removing Redundant Symbols
The intuition behind this notion of redundancy is that 1) if any plan involving the symbol yields a motion sequence that can be rewritten without the symbol, the symbol is redundant, and 2) if we solve a planning instance where a redundant symbol has been removed, we would like to know that the plan is still valid in the original problem. Note that this definition of redundancy is general for any planning domain, although we will use this definition specifically for place symbols that become redundant given the moveRelaxed action. Importantly, removing redundant symbols preserves the feasibility of a planning problem.
@@
Consider a feasible planning instance $R=(\mathcal{P},\mathcal{A},\mathcal{S},\mathcal{O},\mathcal{I}_{0},\mathcal{G})$. For a redundant symbol $x\in\mathcal{O}$, we define a related instance $R^{\prime}=(\mathcal{P},\mathcal{A},\mathcal{S},\mathcal{O}^{\prime},\mathcal{I}_{0}^{\prime},\mathcal{G}^{\prime})$ where $x$ has been removed, i.e., $\mathcal{O}^{\prime}=\mathcal{O}\setminus x$ and $\mathcal{I}_{0}^{\prime}$ contains all facts in $\mathcal{I}_{0}$ except those parameterized by $x$, and similarly for $\mathcal{G}^{\prime}$. Let $\Pi_{R}$ denote the set of valid plans for $R$. Then, $\Pi_{R^{\prime}}\subseteq\Pi_{R}$ and $\Pi_{R^{\prime}}\neq\emptyset$. (Proof deferred to the appendix.)
The requirements for a symbol to be redundant are quite strong (every plan that uses a symbol must have an alternate plan that does not use the symbol and still results in the same motion sequence), but many places in the Inspection Domain have this property given the semantics of moveRelaxed. Removing these places from our task planner’s domain, assuming the motion planner is still aware of them, enables the solver to more efficiently find valid plans that are guaranteed to have also been valid in the un-pruned problem. Moreover, the ability to prune these elements does not restrict the type of goals we are able to specify to our agent, preserving expressivity, while enabling planning at a larger scale.
Proposition 2 (Redundant Places).
Consider a problem instance in the Inspection Domain with no quantifiers that can be parameterized by a place in the goal. A place $p$ is redundant if no facts parameterized by $p$ appear in the initial or goal states, or if (not (VisitedPlace $p$)) appears as a clause in the conjunctive normal form (CNF) of the goal specification. (Proof deferred to appendix.)
We have now identified a potentially large (depending on the sparsity of the goal specification) set of place symbols that can be ignored in the Inspection domain. Our explicit method of defining our problem’s initial state is as follows:
Remark 3.1 (Problem Initialization).
In light of Proposition 2, we only include the following places when instantiating a problem in the Inspection domain: 1) the initial place that the robot is in and 2) any place that appears in the goal. A place $p$ that parameterizes a negated fact (not (VisitedPlace $(\texttt{not }(\texttt{VisitedPlace }p))$ that appears as a clause in the CNF of the goal specification can also be removed.
We have shown that for the Inspection domain, redundant places are very easy to identify and that excluding them from a planning instance adds no computational overhead at runtime. We would like to apply the same idea to similar domains, without needing to reason from scratch about redundancy. We now characterize a sufficient condition on the planning domain structure for places to be redundant. Let $\mathcal{P}_{static}$ denote the set of predicates that do not appear as effects of any action (i.e., they can only be set in the initial state). Let $\mathcal{F}_{static}$ denote the set of facts that correspond to parameterizations of $\mathcal{P}_{static}$. Intuitively, if a domain is structured such that a place can only be parameterized by an action if certain facts hold in the initial state, then it is very easy to check whether a specific place can be used by any actions.
Proposition 1 (Sufficient Conditions for Ignoring Places).
Consider a planning instance $(\mathcal{P},\mathcal{A},\mathcal{S},\mathcal{O},\mathcal{I}_{0},\mathcal{G})$, where for all actions $a_{j}\in\mathcal{A}$ except $a_{j}=\texttt{moveRelaxed}$, satisfying $\text{Pre}(a_{j})$ implies that any place parameterized by $a_{j}$ is in $F_{static}$. In this case, all places that do not parameterize any facts in $\mathcal{F}_{static}$ or the goal are redundant.
For example, if the Inspection domain is augmented with a “report home” action that can only be executed at a designated set of places, then these places (and no others) need to be added to the problem instance. Fig. 2 illustrates how we prune the planning domain.
3.3 Execution Consistency
@@
While our decision to use moveRelaxed to model motion between distance places enables faster planning, it creates a mismatch between the logical and continuous parts of the problem. In the example in Fig. 1B, consider that the robot was also instructed to avoid place $P_{1}$. The ability to include a constraint on the goal states of the form (not (VisitedPlace $P_{1}$)) requires further constraints on the mid- and low-level planners. Executing moveRelaxed from place $P_{0}$ to place $P_{6}$ may involve following a trajectory that takes the robot through place $P_{1}$, even if the goal specifies that $P_{1}$ should not be visited. Technically this is still a valid solution to the planning problem since place $P_{1}$ never appears as a parameter to the moveRelaxed action (and therefore (VisitedPlace B) is not an effect), but clearly the domain with a relaxed movement action does not fully capture the intent of the original planning domain.
@@
To formalize the discrepancy between what happens when the robot executes a motion sequence and the constraints that we expect a planning problem to impose, we introduce the concept of a verifier function. A verifier function maps motion sub-sequences to sets of PDDL domain facts, and “verifies” which additional domain facts would be implicitly true as a result of the agent executing a motion sequence, even if actually adding these facts to the problem instance during the solving process is undesirable computationally. Given a verifier $V$, the facts that hold at each step when executing a motion sequence may be different than expected in the original plan. We denote the facts that would be added by such a verifier applied to the motion sequence associated with $a_{i}$ as $V(a_{i})$, and term this sequence of expanded states the $V$-extended state plan.
@@
For an action plan $\pi=[a_{1},...,a_{n}]$, its corresponding state plan $\mathcal{I}_{\pi}=[\mathcal{I}_{0},...,\mathcal{I}_{n}]$, and verifier function $V$, the V-extended state plan for state $\mathcal{I}_{k}$ is $\mathcal{I}^{\prime}_{1}=\mathcal{I}_{1}\cup V(a_{1})$, and $\mathcal{I}^{\prime}_{k}=\mathcal{I}_{k}\cup\left(\mathcal{I}^{\prime}_{k-1}\setminus\text{Eff}^{-}(a_{k})\right)\cup V(a_{k}).$
The extended state $\mathcal{I}^{\prime}_{k}$ is the state at step $k$ as experienced by the verifier. $\mathcal{I}^{\prime}_{k}$ is composed of the facts $\mathcal{I}_{k}$ in the initial plan, plus any extra facts that were present in the previous extended state $\mathcal{I}^{\prime}_{k-1}$ other than those removed by action $a_{k}$, plus any facts that would be returned by a verifier applied to action $a_{k}$. As discussed in Sec. 2.1, any state plan found by a search algorithm is valid by construction. However, a state plan that is augmented with the extra facts that would be produced by a verifier might not be valid.
Consider a verifier $V_{place}$ that takes a motion sub-sequence $\mu$, and returns a VisitedPlace fact for each place that intersects with the agent’s position while executing $\mu$. For a place $p$ and a trajectory $t$ to be followed by the motion primitive $\texttt{FollowPath}(t)$, we denote $p\cap t$ the section of $t$ that intersects with $p$. We can then define a verifier as
@@
If the motion sequence associated with the action plan would result in the agent visiting a place that we do not expect, then the $V_{place}$-extended state plan would include a VisitedPlace fact that may conflict with the goal. If we care about the robot’s motion respecting the problem’s constraints on visiting certain places, then we need to prove that the $V_{place}$-extended state plan is a valid solution to the planning problem for any instance of the planning domain.
From this idea, we define the concept of execution consistency, which requires that solutions to the planning problem are still valid after considering the facts from a verifier.
Definition 3 (Execution consistent).
@@
A domain is execution consistent with respect to verifier $V$ if, for every valid plan $\pi$, the V-extended state plan is valid.
@@
A domain is trivially execution consistent for the empty verifier $V(\cdot)=\emptyset$, as the extended state plan is equal to the original plan. A domain is also execution consistent if the range of $V$ applied to each motion subsequence $\mu_{i}$ corresponding to action $a_{i}$ is limited to facts in $\text{Eff}(a_{i})$. In other cases, a domain can still be execution consistent for a verifier that would introduce new facts if the domain is carefully crafted. In defining a planning domain for any task, we seek to have it execution consistent with respect to any defined verifiers. If a domain is not execution consistent, then any properties related to predicates in the verifier cannot be guaranteed to hold when executing a plan.